### v2

# Quantum Graph Attention Network: A Novel Quantum Multi-Head Attention Mechanism for Graph Learning

Abstract
We propose the Quantum Graph Attention Network (QGAT), a hybrid graph neural network that integrates variational quantum circuits into the attention mechanism. At its core, QGAT employs strongly entangling quantum circuits with amplitude-encoded node features to enable expressive nonlinear interactions. Distinct from classical multi-head attention that separately computes each head, QGAT leverages a single quantum circuit to simultaneously generate multiple attention coefficients. This quantum parallelism facilitates parameter sharing across heads, substantially reducing computational overhead and model complexity. Classical projection weights and quantum circuit parameters are optimized jointly in an end-to-end manner, ensuring flexible adaptation to learning tasks. Empirical results demonstrate QGAT’s effectiveness in capturing complex structural dependencies and improved generalization in inductive scenarios, highlighting its potential for scalable quantum-enhanced learning across domains such as chemistry, biology, and network analysis. Furthermore, experiments confirm that quantum embedding enhances robustness against feature and structural noise, suggesting advantages in handling real-world noisy data. The modularity of QGAT also ensures straightforward integration into existing architectures, allowing it to easily augment classical attention-based models.
1 Introduction
Graph Neural Networks (GNNs) have emerged as an effective framework for learning from graph-structured data, demonstrating wide applicability in social networks, recommendation systems, bioinformatics, and molecular property prediction[1, 2, 3]. Numerous architectures have been developed to model node dependencies, including Graph Convolutional Networks (GCN)[4, 5], GraphSAGE[6, 7], Graph Isomorphism Networks (GIN)[8], Graph Attention Networks (GAT)[9, 10], GATv2[11], and Graph Transformers[12, 13, 14]. These models typically utilize neighborhood aggregation or attention mechanisms to derive node representations. However, classical GNNs often face representational bottlenecks and scalability issues on large or structurally complex graphs, particularly those with intricate relational patterns or nonlinear interactions[8, 15, 16]. Attention-based models like GAT and GATv2[9, 11, 17] partially address these limitations but still struggle to fully capture dynamic nonlinear interactions within complex graph topologies. Hence, developing more expressive and adaptable architectures is necessary to better exploit the richness inherent in real-world graphs[17].
Recent advancements in quantum computing suggest that quantum machine learning (QML) models may surpass classical methods in capturing complex nonlinear relationships due to fundamental quantum properties, such as entanglement and superposition. These quantum phenomena enable inherently richer representations within high-dimensional Hilbert spaces[18, 19]. Unlike classical neural networks that explicitly introduce nonlinearity through activation functions, quantum models naturally achieve nonlinearity via the unitary evolution of quantum states[20, 21]. For instance, recent studies on quantum pointwise convolution[22] have shown that quantum circuits encode nonlinearities flexibly and scalably, improving feature representations without classical nonlinear activations. Thus, quantum computing holds significant promise for enhancing the expressive capabilities of machine learning models, especially in tasks involving complex data structures challenging traditional approaches.
Quantum attention mechanisms, particularly quantum self-attention, have been demonstrated as theoretically sound and practically feasible on near-term quantum devices. Recent quantum attention models—such as QSAN[23], QKSAN[24], and quantum self-attention networks for text classification[25]—successfully replicate classical self-attention behavior through quantum interference and kernel methods. These models effectively capture long-range dependencies with fewer parameters and improved robustness against overfitting. Furthermore, quantum extensions of Transformer architectures like Quixer[26], Quantum Vision Transformers[27], and QClusformer[28] have shown quantum-enhanced performance across various domains, including image classification, text modeling, and clustering. These findings indicate that quantum attention mechanisms are viable and can outperform classical counterparts in modeling complex data structures, underscoring their potential for integration into graph neural networks.
Motivated by these insights, we propose the Quantum Graph Attention Network (QGAT), a hybrid architecture integrating parameterized quantum circuits within graph attention computations. QGAT aims to enhance the expressive power and robustness of GNNs while preserving the scalability and modularity found in classical models like GAT. Specifically, node features are embedded into quantum states using amplitude encoding, and a single quantum circuit generates multiple attention heads simultaneously. This quantum-inspired design reduces the total parameter count, fosters stronger feature entanglement, and better captures complex relational patterns. Additionally, the intrinsic quantum state distributions contribute to increased robustness against noise, positioning QGAT as an effective and scalable alternative to classical attention-based GNNs.
2 Related Work
Graph Attention Network.
Graph Attention Networks (GAT) [9] introduce a learnable self-attention mechanism into graph neural networks, allowing each node to assign different importance weights to its neighbors during message aggregation. Unlike earlier methods that use fixed or uniform weighting schemes, GAT dynamically computes attention coefficients based on node feature similarity. This enables the model to focus on more relevant neighbors, improving adaptability to heterogeneous graph structures while preserving scalability and parallelizability.
GATv2.
GATv2 [11] improves upon the original GAT by computing attention scores from a joint transformation of both source and target node features before applying the nonlinearity. This change eliminates constraints in the original GAT’s expressiveness, enabling it to capture more complex and asymmetric node relationships. GATv2 retains the computational advantages of GAT while providing strictly greater representational power.
Quantum Attention.
Quantum attention mechanisms aim to enhance classical attention modules by leveraging the high-dimensional, entangled representations offered by quantum systems. Recent studies have shown that variational quantum circuits can function as learnable attention score generators. For example, models such as QSAN [23] and QKSAN [24] use parameterized quantum circuits to produce attention weights, where measured expectation values directly correspond to attention coefficients. These quantum-derived scores inherently capture nonlinear dependencies with fewer parameters. Applications in text classification [25] further demonstrate that quantum self-attention can effectively model semantic relevance, achieving performance on par with or better than classical approaches. These findings support the use of quantum outputs as viable and efficient attention mechanisms.
3 Methods
3.1 Preliminaries
Graph Attention Network.
Graph Attention Networks (GAT) [9] leverage masked self-attention mechanisms to aggregate information from neighboring nodes. Specifically, the attention coefficient $\alpha_{ij}$ between node $i$ and node $j$ is computed as:
$\alpha_{ij}=\frac{\exp\left(\text{LeakyReLU}\left(\mathbf{a}^{\top}[\mathbf{W}\mathbf{h}_{i}\|\mathbf{W}\mathbf{h}_{j}]\right)\right)}{\sum\limits_{k\in\mathcal{N}_{i}}\exp\left(\text{LeakyReLU}\left(\mathbf{a}^{\top}[\mathbf{W}\mathbf{h}_{i}\|\mathbf{W}\mathbf{h}_{k}]\right)\right)},$ (1)
where $\mathbf{W}\in\mathbb{R}^{d^{\prime}\times d}$ is a shared weight matrix applied to input features, $\mathbf{a}\in\mathbb{R}^{2d^{\prime}}$ is a learnable attention vector, $\mathbf{h}_{i},\mathbf{h}_{j}\in\mathbb{R}^{d}$ are input node features, $\mathcal{N}_{i}$ denotes the neighbors of node $i$, and $\cdot\|\cdot$] indicates feature-wise concatenation. This formulation allows each node to assign different attention weights to its neighbors based on their content, thereby enhancing representational flexibility.
GATv2.
GATv2 [11] improves upon the original GAT by allowing the attention mechanism to be input-dependent and more expressive. Instead of computing attention on linearly transformed inputs independently, it applies a joint transformation before nonlinearity:
$\alpha_{ij}=\frac{\exp\left(\mathbf{a}^{\top}\cdot\text{LeakyReLU}(\mathbf{W}[\mathbf{h}_{i}\|\mathbf{h}_{j}])\right)}{\sum\limits_{k\in\mathcal{N}_{i}}\exp\left(\mathbf{a}^{\top}\cdot\text{LeakyReLU}(\mathbf{W}[\mathbf{h}_{i}\|\mathbf{h}_{k}])\right)},$ (2)
where the symbols retain the same meaning as above. By applying the nonlinearity after the joint linear transformation of both source and target node features, GATv2 eliminates the constraint of static attention patterns and enables asymmetric relational modeling between nodes. This makes GATv2 strictly more expressive while maintaining the efficiency and parallelism of the original design.
Quantum Machine Learning.
Quantum Machine Learning (QML) integrates quantum computation into machine learning pipelines by leveraging quantum circuits to process and transform data. A core step in QML is quantum data embedding, which encodes classical input vectors into quantum states suitable for further quantum processing. Two widely used strategies are angle embedding and amplitude embedding.
In angle embedding, each classical feature is mapped to a rotation gate acting on a corresponding qubit:
$|\psi(\mathbf{x})\rangle=\bigotimes_{i=1}^{n}R_{Y}(x_{i})|0\rangle,$ (3)
where $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})$ is the input vector and $R_{Y}(x_{i})=\exp(-ix_{i}Y/2)$ denotes a rotation around the Y-axis on the $i$-th qubit. This method creates shallow, hardware-friendly circuits for encoding real-valued features.
In contrast, amplitude embedding directly maps the components of $\mathbf{x}$ into the amplitude vector of a quantum state:
$|\psi(\mathbf{x})\rangle=\sum_{i=0}^{2^{n}-1}x_{i}|i\rangle,$ (4)
where $\mathbf{x}$ is normalized such that $\sum_{i}|x_{i}|^{2}=1$. This yields an exponentially compact representation of high-dimensional data [29, 30], but often demands deep and complex state preparation circuits.
After embedding, the quantum state is processed through a parameterized quantum circuit, commonly referred to as a variational quantum circuit (VQC). These circuits apply trainable unitary operations and are optimized using classical algorithms such as gradient descent or Adam to minimize task-specific loss functions, including mean squared error or cross-entropy [31]. The expressibility of the variational ansatz, together with the loss landscape and optimizer choice, plays a critical role in convergence and final model performance [31, 32].
3.2 Quantum Graph Attention Network
Motivation and Overview.
The Quantum Graph Attention Network (QGAT) integrates quantum circuits into the attention mechanism to enhance feature representation. Unlike classical models such as GAT and GATv2, which rely on linear or piecewise-linear transformations, QGAT leverages the nonlinear and high-dimensional mappings of quantum circuits induced by unitary evolution and entanglement. This enables richer feature interactions with fewer parameters. By embedding quantum layers into attention computation, QGAT achieves greater expressiveness without sacrificing scalability.
Quantum Data Embedding.
QGAT adopts amplitude encoding to map node features into quantum states. Given a classical input vector $\mathbf{x}\in\mathbb{R}^{d}$, it is first normalized and padded to length $2^{n}$ (if necessary), and then encoded as:
$|\psi(\mathbf{x})\rangle=\sum_{i=0}^{2^{n}-1}x_{i}|i\rangle,$ (5)
where $n=\lceil\log_{2}d\rceil$ is the number of qubits. This approach enables compact and expressive representation of high-dimensional node features using a logarithmic number of qubits.
Amplitude encoding is chosen for its ability to preserve global feature structure by directly mapping inputs to quantum state amplitudes. Compared to angle-based methods, it captures richer feature relationships with fewer gates and requires only logarithmic qubits relative to input size, reducing hardware demands while maintaining expressiveness.
Quantum Circuit Design.
The quantum module in QGAT adopts a variant of variational quantum circuits based on strongly entangling layers [33]. Each block begins with an amplitude-encoded input and applies parameterized single-qubit rotations followed by entangling operations.
Each single-qubit gate is implemented using a standard Z-Y-Z decomposition:
$G_{j}=R_{z}(\mu_{1})R_{y}(\mu_{2})R_{z}(\mu_{3}),$ (6)
where $\mu_{1}$, $\mu_{2}$, and $\mu_{3}$ are trainable parameters, and $R_{y}(\theta)$, $R_{z}(\theta)$ denote Y- and Z-axis rotations, respectively. This decomposition is both universal and hardware-friendly, allowing expressive and stable control over state evolution with minimal circuit depth.
Entanglement is introduced using two-qubit gates placed after the single-qubit rotations. We use imprimitive two-qubit operations capable of entangling initially separable states. A representative example is the controlled single-qubit gate:
$C_{a}(G_{b}):|x\rangle|y\rangle\mapsto|x\rangle\otimes G_{b}^{x}|y\rangle,$ (7)
where $G$ is a single-qubit gate applied to qubit $b$ conditioned on the state of qubit $a$. In the special case $G=X$, this becomes a standard CNOT gate.
The qubit connectivity for entanglement is governed by a tunable range parameter, alternating across layers to allow both short- and long-range entanglement. A complete quantum block consists of a sequence of single-qubit rotations and controlled entangling gates:
$B=\prod_{k=0}^{n-1}R^{Y}_{k}\,C_{ck}(P_{tk})\,\prod_{j=0}^{n-1}G_{j},$ (8)
where $R^{Y}_{k}$ denotes a Y-rotation on qubit $k$, $C_{ck}(P_{tk})$ is a controlled phase gate, and $G_{j}$ is the Z-Y-Z rotation gate described above.
By stacking multiple such blocks, the circuit encodes classical features into high-dimensional entangled quantum states. This enhances the attention mechanism’s capacity to model complex interactions. An example with five qubits and two entangling blocks is shown in Figure 1, illustrating how entanglement is distributed across the entire system.
Figure 1: Example of a quantum circuit used in QGAT with $n=5$ qubits and two strongly entangling blocks. Each block includes parameterized single-qubit rotations ($R_{y}$, $R_{z}$) and entangling layers constructed using CNOT gates. After the final block, Pauli-$Z$ measurements are performed. The two-qubit entangling gates are CNOTs applied chronologically on qubits $i=1,\dots,M$, where the target qubit is $(i+r)\bmod M$ and $r$ is a tunable range parameter with $0<r<M$. This configuration allows the circuit to achieve both local and long-range entanglement.
Quantum Graph Attention Architecture.
The proposed Quantum Graph Attention mechanism computes attention coefficients using a variational quantum circuit. Given input node features $\mathbf{h}\in\mathbb{R}^{N\times d}$, we first project them into multi-head representations via a shared linear transformation $\mathbf{W}\in\mathbb{R}^{d\times(h\cdot d_{\text{out}})}$. For each edge $(i,j)$, the projected features are concatenated:
$\mathbf{a}_{ij}=[\mathbf{W}\mathbf{h}_{i}\|\mathbf{W}\mathbf{h}_{j}]\in\mathbb{R}^{2hd_{\text{out}}},$ (9)
where $\cdot\|\cdot$] denotes feature-wise concatenation.
To construct the quantum input, we further concatenate the original node features $\mathbf{h}_{i}$ and $\mathbf{h}_{j}$ with $\mathbf{a}_{ij}$ and apply a second projection $\mathbf{P}$:
$\mathbf{a}^{\prime}_{ij}=\mathbf{P}([\mathbf{a}_{ij}\|\mathbf{h}_{i}\|\mathbf{h}_{j}])\in\mathbb{R}^{2^{n_{q}}\cdot\lceil h/n_{q}\rceil},$ (10)
where $n_{q}$ is the number of qubits in the quantum circuit.
To preserve essential structural and semantic information, we introduce a residual-enhanced design that concatenates the original node features with projected multi-head features. This enriches the quantum encoding and mitigates information loss from earlier transformations. The projection matrix $\mathbf{P}$ adjusts the input length to match the amplitude encoding dimension while retaining key feature correlations. The resulting vector $\mathbf{a}^{\prime}_{ij}$ is then amplitude-encoded and processed by a quantum circuit. Although $\mathbf{P}$ compresses the input, the residual path helps retain fine-grained details. The circuit outputs multiple attention logits by measuring Pauli-$Z$ expectation values across qubits. If more heads are needed than available qubits, the circuit is executed repeatedly with the same input.
The quantum circuit applies the variational ansatz $U(\boldsymbol{\theta})$ to the amplitude-encoded state $|\psi(\mathbf{a}^{\prime}_{ij})\rangle$. Each qubit $k$ produces one attention logit by measuring the expectation value of the Pauli-$Z$ observable:
$e_{ij}^{(k)}=\langle\psi(\mathbf{a}^{\prime}_{ij})|U^{\dagger}(\boldsymbol{\theta})Z_{k}U(\boldsymbol{\theta})|\psi(\mathbf{a}^{\prime}_{ij})\rangle,\quad k=1,\dots,n_{q},$ (11)
where $Z_{k}$ is the Pauli-$Z$ operator acting on the $k$-th qubit.
When the number of required attention heads $h$ exceeds the available qubits $n_{q}$, the circuit is executed $n_{\mathrm{exec}}$ times:
$n_{\mathrm{exec}}=\left\lceil\frac{h}{n_{q}}\right\rceil,$ (12)
and the outputs from all executions are concatenated to form the full set of $h$ attention logits.
These logits are normalized using softmax:
$\alpha_{ij}^{(k)}=\frac{\exp(e_{ij}^{(k)})}{\sum\limits_{m\in\mathcal{N}_{i}}\exp(e_{im}^{(k)})},$ (13)
where $\mathcal{N}_{i}$ denotes the neighbors of node $i$, and $m$ iterates over all such neighbors in the normalization term.
No additional nonlinearities (e.g., LeakyReLU) are applied to $e_{ij}^{(k)}$, since the quantum circuit itself induces nonlinearity through its parameterized unitaries and measurement process.
The updated node representation is aggregated as follows:
Concat:
$\mathbf{h}^{\prime}i=\sigma\left(\mathrm{concat}{k=1}^{h}\sum_{j\in\mathcal{N}i}\alpha{ij}^{(k)}\mathbf{W}^{(k)}\mathbf{h}_{j}\right),$ (14)
Mean:
$\mathbf{h}^{\prime}i=\sigma\left(\frac{1}{h}\sum{k=1}^{h}\sum_{j\in\mathcal{N}i}\alpha{ij}^{(k)}\mathbf{W}^{(k)}\mathbf{h}_{j}\right),$ (15)
where $\mathbf{W}^{(k)}$ denotes the value projection matrix for the $k$-th head, and $\sigma$ is a non-linear activation function.
Input: Node features $\mathbf{h}\in\mathbb{R}^{N\times d}$, edge index $\mathcal{E}$, number of attention heads $h$, number of qubits $n_{q}$, number of quantum layers $L$, classical projection weights $\mathbf{W}$, compression matrix $\mathbf{P}$, activation $\sigma$, merge type (concat or mean).
Output: Updated node features $\mathbf{h}^{\prime}\in\mathbb{R}^{N\times d^{\prime}}$.
1
2Project node features into multi-head space: $\mathbf{h}^{\prime}_{i}\leftarrow\mathbf{W}\mathbf{h}_{i}$, for all $i$
3foreach edge $(i,j)\in\mathcal{E}$ do
4   Concatenate projected and raw features: $\mathbf{a}_{ij}\leftarrow[\mathbf{h}^{\prime}_{i}\|\mathbf{h}^{\prime}_{j}\|\mathbf{h}_{i}\|\mathbf{h}_{j}]$
5  Compress to match amplitude encoding input: $\mathbf{a}^{\prime}_{ij}\leftarrow\mathbf{P}(\mathbf{a}_{ij})\in\mathbb{R}^{2^{n_{q}}\cdot\lceil h/n_{q}\rceil}$
6  Initialize quantum parameters $\boldsymbol{\theta}$
7 for $m\leftarrow 1$ to $\lceil h/n_{q}\rceil$ do
8      Amplitude encode $\mathbf{a}^{\prime}_{ij}$ as $|\psi(\mathbf{a}^{\prime}_{ij})\rangle$
9      Apply variational circuit $U(\boldsymbol{\theta})$ with strongly entangling layers: $|\psi^{\prime}_{ij}\rangle\leftarrow U(\boldsymbol{\theta})|\psi(\mathbf{a}^{\prime}_{ij})\rangle$
10      Measure Pauli-$Z$ expectation on all $n_{q}$ qubits: $e_{ij}^{(k)}\leftarrow\langle\psi(\mathbf{a}^{\prime}_{ij})|U^{\dagger}Z_{k}U|\psi(\mathbf{a}^{\prime}_{ij})\rangle$
11
12
13Normalize logits via softmax: $\alpha_{ij}^{(k)}\leftarrow\frac{\exp(e_{ij}^{(k)})}{\sum\limits_{m\in\mathcal{N}_{i}}\exp(e_{im}^{(k)})}$
14Update node embeddings:
15 if merge = concat then
16 $\mathbf{h}^{\prime}_{i}\leftarrow\sigma\left(\mathrm{concat}_{k=1}^{h}\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}^{(k)}\mathbf{W}^{(k)}\mathbf{h}_{j}\right)$
17else if merge = mean then
18 $\mathbf{h}^{\prime}_{i}\leftarrow\sigma\left(\frac{1}{h}\sum_{k=1}^{h}\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}^{(k)}\mathbf{W}^{(k)}\mathbf{h}_{j}\right)$
19
20Apply residual connection and dropout
return updated node embeddings $\mathbf{h}^{\prime}$
Algorithm 1 Quantum Graph Attention Network (QGAT) Layer
4 Evaluation
We evaluate the performance of QGAT through a series of experiments aimed at assessing its robustness, expressiveness, and generalization across multiple graph learning tasks. Section 4.1 examines the model’s resilience to feature and structural perturbations. Section 4.2 presents benchmark comparisons for node-level tasks, while Section 4.3 evaluates link prediction performance. Across all experiments, we compare QGAT with established baselines, including GAT and GATv2, to highlight the benefits of quantum-enhanced attention. The evaluations cover both transductive and inductive settings to assess generalization to unseen nodes and graphs. Full training configurations—covering optimizer settings, learning rates, batch sizes, and early stopping criteria—are provided in Appendix A. All reported results reflect test set performance, evaluated at the best validation checkpoint over five independent runs.
4.1 Robustness to Noise
We evaluate the robustness of QGAT on the ogbn-arxiv dataset, a widely used citation graph benchmark with rich node features and semantic labels. This dataset is selected for its moderate size, structural complexity, and frequent use in evaluating graph learning models. We assess the model’s stability under two types of perturbations: feature noise and structural noise.
Feature Noise.
To test robustness against feature corruption, we inject Gaussian noise into the node features and compare QGAT’s performance to that of GAT and GATv2. Because QGAT uses amplitude encoding—which directly maps features into quantum state amplitudes—it is particularly important to assess its sensitivity to such distortions.
For a given noise level $\epsilon$, each input feature vector $\mathbf{x}$ is perturbed by adding zero-mean Gaussian noise:
$\tilde{\mathbf{x}}=\mathbf{x}+\epsilon\cdot\mathcal{N}(0,\mathbf{I}),$ (16)
where $\epsilon\in{0.0,0.01,0.05,0.1,0.2}$ controls the magnitude of the perturbation. The same noise level is applied during both training and evaluation to ensure consistency.
Figure 2(a) shows test accuracy under increasing noise levels. While all models degrade with stronger perturbations, QGAT consistently outperforms GAT and GATv2 across all levels. This robustness is attributed to the global nature of amplitude embedding and the expressiveness introduced by quantum entanglement, which together help QGAT retain discriminative information even in noisy conditions.
Structural Noise.
To assess QGAT’s robustness to topological perturbations, we introduce structural noise by randomly adding edges to the graph. We evaluate a range of predefined noise ratios $\eta\in{0.0,0.1,0.2,0.3,0.4,0.5}$, where each value denotes the proportion of additional edges relative to the original edge count.
Given an undirected graph with $E$ edges and NN nodes, we sample $\lfloor\eta\cdot E\rfloor$ new edges by selecting random node pairs $(u,v)$ such that $u\neq v$ and $(u,v)\notin\mathcal{E}$. These randomly generated edges form a perturbation set $\mathcal{E}{\text{rand}}$, which is added to the original edge index:
$\mathcal{E}{\text{noisy}}=\mathcal{E}\cup\mathcal{E}_{\text{rand}}.$ (17)
Figure 2(b) presents the test accuracy under increasing structural noise levels. As noise increases, the performance of all models declines. However, QGAT consistently outperforms both GAT and GATv2, maintaining a higher accuracy margin especially under moderate to severe noise. This result suggests that QGAT is more resilient to irregular or corrupted graph topologies, likely due to the expressive relational modeling induced by amplitude encoding and quantum entanglement.
(a) Feature noise. Test accuracy under Gaussian perturbation with $\sigma$.
(b) Structural noise. Test accuracy under random edge insertion with $\sigma$.
Figure 2: Robustness comparison of QGAT, GATv2, and GAT under feature and structural noise. Each point is averaged over 5 runs with standard deviation shown as error bars.
4.2 Node Property Prediction
We evaluate the effectiveness of QGAT in node property prediction tasks under two settings: transductive, where the entire graph is available during training, and inductive, where the model must generalize to unseen nodes or subgraphs. These settings offer complementary insights into the model’s ability to capture both local neighborhood patterns and broader structural dependencies.
Transductive
We conduct experiments on three widely used transductive node classification benchmarks: Pubmed, ogbn-arxiv, and ogbn-products. These datasets test the model’s capacity to learn from full graph structures, with access to all nodes and edges during training. QGAT is evaluated under the same experimental settings as classical baselines, including GAT and GATv2. As shown in Table 1, QGAT consistently achieves higher accuracy across all datasets, demonstrating its superior ability to capture structural and semantic node-level relationships.
Table 1: Accuracy (%) on transductive node classification benchmarks. All results are averaged over 5 runs; standard deviations are reported.
Model Pubmed ogbn-arxiv ogbn-products
GAT [11] $78.1\pm 0.59$ $71.54\pm 0.3$ $79.04\pm 1.54$
GATv2 [11] $78.5\pm 0.38$ $71.87\pm 0.25$ $80.63\pm 0.7$
QGAT $\mathbf{79.2\pm 0.62}$ $\mathbf{73.62\pm 0.42}$ $\mathbf{82.10\pm 2.31}$
Based on the results in Table 1, QGAT achieves the highest accuracy across all three transductive node classification benchmarks. It surpasses GATv2 by approximately $0.7\%$ on Pubmed, $1.75\%$ on ogbn-arxiv, and $1.47\%$ on ogbn-products. These consistent improvements, accompanied by low standard deviations, indicate that QGAT offers a more expressive and stable mechanism for capturing node-level dependencies when the full graph is available during training.
Inductive.
To evaluate the generalization ability of QGAT under inductive settings, we conduct node classification benchmarks on two widely used datasets: PPI and ogbn-proteins. These datasets involve either multiple disjoint graphs (PPI) or large-scale protein interaction networks (ogbn-proteins), where test nodes or subgraphs are inaccessible during training. This setup simulates realistic scenarios where models must transfer learned representations to previously unseen structures. As summarized in Table 2, QGAT is evaluated alongside GAT and GATv2 under identical configurations.
Table 2: Inductive node classification performance on PPI and ogbn-proteins. Metric: Micro-F1 (PPI) and ROC-AUC (ogbn-proteins). All results are averaged over 5 runs; standard deviations are reported.
Model PPI (Micro-F1) ogbn-proteins (ROC-AUC)
GAT [6] $97.3\pm 0.20$ $78.63\pm 1.62$
GATv2 $98.2\pm 0.25$ $\mathbf{79.52\pm 0.55}$ [11]
QGAT $\mathbf{98.9\pm 0.12}$ $79.41\pm 0.21$
QGAT demonstrates strong inductive generalization across both datasets. On PPI, it outperforms both GAT and GATv2, achieving a Micro-F1 score of $98.9\%$, reflecting its ability to transfer representations across disconnected graphs. On ogbn-proteins, QGAT attains a ROC-AUC of $79.41\%$, performing competitively with GATv2 while exhibiting lower variance, indicating more stable prediction across protein interaction structures.
Since prior literature does not report GATv2 results on the PPI dataset, we implemented a three-layer GATv2Conv [11] baseline using attention heads $[8,8,4]$, hidden dimension of 256256, dropout rate of $0.5$, and LayerNorm applied after each layer.
4.3 Link Prediction
To assess QGAT’s ability to capture implicit relationships between node pairs, we evaluate its performance on link prediction tasks using two representative datasets: ogbl-collab and ogbl-citation2. These datasets span diverse domains including collaboration networks, drug-drug interactions, and citation graphs. The model is trained to predict the existence of edges based on learned node representations, providing insight into its capacity for structural reasoning and generalization.
Table 3: Link prediction performance on OGB benchmarks. Metric: Hits@50 (%) for ogbl-collab, and Mean Reciprocal Rank (MRR) for ogbl-citation2. All results are averaged over 5 runs; standard deviations are reported.
Model ogbl-collab ogbl-citation2
GAT [11] $46.63\pm 2.80$ $75.95\pm 1.31$
GATv2 [11] $49.7\pm 3.08$ $80.14\pm 0.71$
QGAT $\mathbf{51.2\pm 1.92}$ $\mathbf{82.2\pm 1.27}$
As shown in Table 3, QGAT achieves the highest Hits@50 on ogbl-collab with $51.2\%$, outperforming GAT and GATv2 by $4.57\%$ and $1.5\%$, respectively. On the more challenging ogbl-citation2 benchmark, QGAT attains the top mean reciprocal rank (MRR) of $82.2\%$, surpassing both baselines. These results highlight QGAT’s enhanced ability to model latent structural dependencies, driven by the nonlinear and entangled transformations introduced through the quantum attention mechanism.
5 Conclusion
We introduced the Quantum Graph Attention Network (QGAT), a hybrid architecture integrating variational quantum circuits into graph neural network attention mechanisms. By using amplitude encoding and strongly entangled quantum layers, QGAT captures rich nonlinear interactions and efficiently implements multi-head attention within a single quantum circuit. This design reduces parameter overhead and enhances representational capacity, enabling compact yet expressive models. Experiments across various benchmarks—including transductive and inductive node classification, as well as link prediction—show that QGAT consistently outperforms classical baselines such as GAT and GATv2. Notably, these performance gains persist even under strict hyperparameter constraints due to memory limitations, underscoring the model’s efficiency and robust generalization. Moreover, evaluations under noisy conditions highlight the increased robustness of amplitude-encoded quantum representations against feature-level and structural perturbations. Thus, QGAT effectively demonstrates the practicality and benefits of quantum-enhanced attention in realistic graph learning tasks.
Discussion and Future Work.
Our experiments demonstrate that QGAT excels on complex, multi-label datasets such as PPI, highlighting its effectiveness in capturing intricate relational structures found in biological interaction networks and molecular systems. Additionally, QGAT shows competitive results on link prediction tasks, reflecting its capacity to learn fine-grained edge-level relationships. Thanks to its modular design, QGAT can directly replace classical GAT or GATv2 layers, facilitating easy integration into various existing architectures and training workflows.
However, hardware scalability remains a key limitation. Our current implementation does not support distributed quantum-classical execution across multiple GPUs or nodes, restricting hyperparameter configurations and limiting applicability to large-scale graphs. Addressing this constraint—through circuit batching, tensor parallelism, or quantum-aware distributed scheduling—is essential for scaling QGAT to larger and more computationally intensive graph tasks in the future.
Appendix A Training Details
QGAT is implemented with Pennylane[35] for quantum circuit simulation and PyTorch for classical computation. All experiments are conducted on a single NVIDIA H100 GPU.
The quantum module introduces additional memory overhead due to circuit simulation and multi-head execution. As a result, we restrict the hyperparameter search space and do not explore larger hidden dimensions or wider attention heads. Despite this limitation, QGAT consistently achieves competitive performance across all benchmarks.
A.1 Robustness to Noise
For QGAT, we perform a grid search over the following hyperparameters: learning rate $\in\{1\times 10^{-2},2\times 10^{-3},1\times 10^{-3},1\times 10^{-4}\}$, hidden dimension $\in\{32,64,128\}$, and dropout rate $\in\{0.4,0.5,0.6\}$, and strong entanglement layer $\in\{2,3,4\}$. The model uses three graph attention layers with attention heads configured as $[4,4,4]$ or $[6,6,4]$, depending on the dataset. Optimization is performed using the AdamW optimizer along with a CosineAnnealingLR scheduler. All training is carried out using mixed-precision (AMP) to reduce memory usage and accelerate convergence.
Experiments on the ogbn-arxiv dataset are conducted using the GraphSAINT mini-batch sampling strategy [36] to reduce memory consumption and enable scalable training. This sampling approach allows QGAT to operate effectively on large graphs without exceeding GPU memory limits.
For baseline models (GAT and GATv2), we adopt the hyperparameter configurations reported in the original papers [9, 11] to ensure fair and consistent comparisons.
A.2 Node Property Prediction
Transductive
We evaluate QGAT on three node classification benchmarks: Pubmed, ogbn-arxiv, and ogbn-products, each with dataset-specific hyperparameter configurations. Across all datasets, training is performed using the AdamW optimizer with a CosineAnnealingLR scheduler. Mixed-precision training (AMP) is employed to reduce GPU memory usage and improve efficiency.
For the Pubmed dataset, we perform a grid search over the following hyperparameters: learning rate $\in\{1\times 10^{-2},2\times 10^{-3},1\times 10^{-3},1\times 10^{-4}\}$, hidden dimension $\in\{8,16,32\}$, and dropout $\in\{0.3,0.4,0.5,0.6\}$, and strong entanglement layer $\in\{2,3,4\}$. The architecture uses two QGAT layers, with attention head settings including $[6,4]$, $[4,1]$, and $[4,4]$.
For the ogbn-arxiv dataset, we adopt the same hyperparameter configurations as in the robustness experiments: learning rate $\in\{1\times 10^{-2},2\times 10^{-3},1\times 10^{-3},1\times 10^{-4}\}$, hidden dimension $\in\{32,64,128\}$, dropout $\in\{0.4,0.5,0.6\}$, three layers, and attention head configurations $[4,4,4]$ or $[6,6,4]$, and strong entanglement layer $\in\{2,3,4\}$. Training on ogbn-arxiv is performed with the GraphSAINT sampling strategy to enable scalable training on the full graph.
For the ogbn-products dataset, we use a similar setup to ogbn-arxiv, with a grid over learning rate $\in\{1\times 10^{-2},2\times 10^{-3},1\times 10^{-3},1\times 10^{-4}\}$, hidden dimension $\in\{32,64,128\}$, dropout $\in\{0.4,0.5,0.6\}$, and attention heads $[4,4,4]$ or $[6,6,4]$ across three QGAT layers, and strong entanglement layer $\in\{2,3,4\}$. Due to the scale of the dataset, GraphSAINT is also used here to reduce computational and memory overhead.
Inductive.
For the PPI dataset, we perform a grid search over learning rate $\in\{1\times 10^{-2},2\times 10^{-3},1\times 10^{-3},1\times 10^{-4}\}$, hidden dimension $\in\{32,64,128,256\}$, and dropout $\in\{0.0,0.3,0.5\}$, and strong entanglement layer $\in\{2,3,4\}$. The model consists of three QGAT layers with attention head configurations $[6,6,4]$ and $[8,8,4]$. Optimization is carried out using AdamW with a CosineAnnealingLR scheduler. All training is performed using mixed-precision training (AMP), and a fixed batch size of 2 is used for training, validation, and testing to accommodate GPU memory constraints.
For the ogbn-proteins dataset, we use a similar setup, with learning rate $\in\{1\times 10^{-2},2\times 10^{-3},1\times 10^{-3},1\times 10^{-4}\}$, hidden dimension $\in\{32,64,128\}$, dropout $\in\{0.0,0.3,0.5\}$, three QGAT layers using either $[6,6,4]$ or $[8,8,4]$ heads, and strong entanglement layer $\in\{2,3,4\}$. GraphSAINT is applied for scalable mini-batch training due to the dataset’s large size. The model is optimized with AdamW and a CosineAnnealingLR scheduler, using mixed-precision training to reduce memory usage and improve runtime efficiency.
A.3 Link Prediction
For the ogbl-collab dataset, we perform a grid search over learning rate $\in\{1\times 10^{-2},2\times 10^{-3},1\times 10^{-3},1\times 10^{-4}\}$, hidden dimension $\in\{64,128,256\}$, and dropout rate $\in\{0.3,0.4,0.5,0.6\}$, and strong entanglement layer $\in\{2,3,4\}$. The model is composed of three QGAT layers with attention heads configured as $[4,4,4]$ or $[6,6,4]$. To handle the large scale of the dataset, we apply GraphSAINT [36] for sampling-based training. Optimization is done using AdamW with a CosineAnnealingLR scheduler, and mixed-precision training is used throughout.
For the ogbl-citation2 dataset, we adopt a similar setup with adjusted learning rate grid $\in\{2\times 10^{-3},1\times 10^{-3},1\times 10^{-4},5\times 10^{-5}\}$, while using the same hidden dimension range $\{64,128,256\}$, dropout values $\{0.3,0.4,0.5,0.6\}$, strong entanglement layer $\in\{2,3,4\}$, and three-layer QGAT architecture. Attention heads are also set to $[4,4,4]$ or $[6,6,4]$, and GraphSAINT is employed to enable efficient training. The optimizer and scheduler choices remain consistent with previous settings.
Appendix B Dataset details
B.1 Node Property Prediction
This section summarizes the datasets used for node property prediction tasks, including both transductive and inductive settings. Table 4 reports key statistics such as the number of nodes, edges, feature dimensions, and task type. These datasets span various domains and settings, providing a diverse benchmark to evaluate the generalization and robustness of QGAT. In particular, Pubmed, ogbn-arxiv, and ogbn-products are evaluated under transductive learning, while PPI and ogbn-proteins are used in inductive scenarios.
Table 4: Statistics and settings of datasets used for node property prediction.
Dataset #Nodes #Edges #Features Task
Pubmed 19,717 44,338 500 Node classification
ogbn-arxiv 169,343 1,166,243 128 Node classification
ogbn-products 2,449,029 61,859,140 100 Node classification
PPI 56,944 818,716 50 Node classification (multi-label)
ogbn-proteins 132,534 39,561,252 8 (edge features) Node classification (binary)
B.2 Link Prediction
We use two benchmark datasets from the Open Graph Benchmark (OGB) suite to evaluate the link prediction capability of QGAT. These datasets span different domains—collaboration networks and citation graphs—challenging the model to capture implicit structural relationships between node pairs. Table 5 summarizes the key statistics and task definitions.
Table 5: Statistics and settings of datasets used for link prediction.
Dataset #Nodes #Edges #Features Task
ogbl-collab 235,868 1,285,465 128 Link prediction (Hits@50)
ogbl-citation2 2,927,963 30,000,000 128 Link prediction (MRR)
Appendix C Parameter Comparison
Table 6: Parameter comparison of GAT, GATv2, and QGAT on the PPI dataset.
Model Hidden Dim Layers Attention Heads Parameters
GAT 256 3 [8, 8, 4] 6,439,545
GATv2 256 3 [8, 8, 4] 12,838,521
QGAT (Strong Entanglement Layer = 2) 256 3 [8, 8, 4] 6,720,017
QGAT (Strong Entanglement Layer = 3) 256 3 [8, 8, 4] 6,720,053
QGAT (Strong Entanglement Layer = 4) 256 3 [8, 8, 4] 6,720,089
Table 6 shows that QGAT, despite integrating quantum circuits and employing strong entanglement, has a parameter count only slightly higher than GAT and significantly lower than GATv2. This indicates that QGAT effectively enhances nonlinear expressivity and feature interactions with minimal additional parameters, achieving a balance between performance and parameter efficiency.
However, due to the use of a PennyLane quantum simulator combined with PyTorch in hybrid mode, training efficiency has not been fully optimized. As a result, the training time of QGAT is approximately five to six times longer than that of GAT and GATv2 under the same hardware conditions.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Yes, the abstract and introduction clearly describe the main contributions, including the design of QGAT and its performance improvements over classical baselines. These claims are supported by both the model description and experimental results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper includes a discussion of QGAT’s limitations, particularly in terms of computational cost and scalability due to quantum circuit simulation, as well as the restricted hyperparameter space caused by GPU memory constraints.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not present formal theoretical results that require assumptions or proofs. The contributions are primarily empirical and architectural, focusing on model design, implementation, and evaluation.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper provides detailed architectural descriptions, training hyperparameters, and experimental configurations for all tasks. This includes dataset-specific training details, model variants, optimizer and scheduler choices, and evaluation metrics. These details are sufficient to reproduce the main experimental results, even without code release.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: The code is currently not released due to ongoing downstream application development. We plan to make it available after completing further internal experiments.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper specifies all key training and evaluation settings, including dataset splits, optimizer, learning rate schedules, dropout rates, number of layers and attention heads, and other hyperparameters. Detailed training configurations for each benchmark task (transductive, inductive, and link prediction) are provided in the Appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: All experimental results are reported with standard deviation error bars calculated over five independent runs using different random seeds. The error bars in plots and tables reflect variability due to training initialization and stochastic optimization. We state explicitly that the bars represent 1-sigma (standard deviation) and reference the corresponding figures (e.g., Figure 2(a), Figure 2(b)) and tables (e.g., Table 1, Table 2) in the main text.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: All experiments were conducted using a single NVIDIA H100 GPU with 80GB memory. Training time per run varied by dataset, ranging from approximately 0.5–3 hours for node classification tasks to 3–6 hours for large-scale link prediction tasks. Memory usage was managed using mixed-precision training and sampling strategies (e.g., GraphSAINT).
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research adheres to the NeurIPS Code of Ethics. It involves no human subjects, personally identifiable information, sensitive content, or dual-use concerns. All datasets used are publicly available and widely adopted in graph learning research.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: This work explores a quantum-enhanced graph attention architecture (QGAT) for graph learning tasks. Potential positive societal impacts include improved performance in scientific domains involving complex graph data, such as drug discovery, protein interaction prediction, and molecular property analysis. These applications may accelerate biomedical research and material science.
    However, as with most machine learning models, the method could potentially be misused in settings like surveillance or social network profiling, where enhanced modeling of graph structures could amplify privacy risks. While our work focuses on foundational research and does not target specific applications, we acknowledge this dual-use nature and emphasize that ethical deployment requires domain-specific safeguards and oversight.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper does not involve the release of high-risk models or scraped datasets. QGAT is a foundational architecture evaluated on publicly available benchmark datasets (e.g., OGB, PPI), and does not introduce or train generative models, language models, or datasets with potential safety concerns. Therefore, specific safeguards are not applicable in this context.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All datasets and software frameworks used in this work are publicly available under appropriate open licenses. The datasets (e.g., OGB datasets, PPI) are accessed via the Open Graph Benchmark (OGB) platform, which is released under the MIT license. The quantum simulation backend is implemented using Pennylane, which is licensed under Apache 2.0, and the classical components are built with PyTorch, licensed under BSD. Citations and references are provided for all external assets used.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., $C$-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This paper does not introduce any new datasets, models, or code assets. All experiments are conducted on publicly available benchmarks and existing libraries, with no new data collection or external asset release.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve any experiments with human subjects or the use of crowdsourcing platforms. All evaluations are conducted on publicly available benchmark datasets using automated computational methods.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification:This work does not involve any research with human subjects or data collected from human participants. All evaluations are performed on publicly available benchmark datasets using computational methods without human intervention.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: This paper does not incorporate LLMs as part of the core methodology, nor are LLMs used in any original or non-standard way that impacts the scientific contributions of the research. All LLM usage, if any, was limited to language editing and does not affect the technical content.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v3

@@
 We evaluate a range of predefined noise ratios $\eta\in{0.0,0.1,0.2,0.3,0.4,0.5}$, where each value denotes the proportion of additional edges relative to the original edge count.
-Given an undirected graph with $E$ edges and NN nodes, we sample $\lfloor\eta\cdot E\rfloor$ new edges by selecting random node pairs $(u,v)$ such that $u\neq v$ and $(u,v)\notin\mathcal{E}$.
+Given an undirected graph with $E$ edges and $N$ nodes, we sample $\lfloor\eta\cdot E\rfloor$ new edges by selecting random node pairs $(u,v)$ such that $u\neq v$ and $(u,v)\notin\mathcal{E}$.
 These randomly generated edges form a perturbation set $\mathcal{E}{\text{rand}}$, which is added to the original edge index:
@@
 On ogbn-proteins, QGAT attains a ROC-AUC of $79.41\%$, performing competitively with GATv2 while exhibiting lower variance, indicating more stable prediction across protein interaction structures.
-Since prior literature does not report GATv2 results on the PPI dataset, we implemented a three-layer GATv2Conv [11] baseline using attention heads $[8,8,4]$, hidden dimension of 256256, dropout rate of $0.5$, and LayerNorm applied after each layer.
+Since prior literature does not report GATv2 results on the PPI dataset, we implemented a three-layer GATv2Conv [11] baseline using attention heads $[8,8,4]$, hidden dimension of $256$, dropout rate of $0.5$, and LayerNorm applied after each layer.
 4.3 Link Prediction
@@
 As a result, the training time of QGAT is approximately five to six times longer than that of GAT and GATv2 under the same hardware conditions.
-NeurIPS Paper Checklist
-The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact.
-Do not remove the checklist: The papers not including the checklist will be desk rejected.
-The checklist should follow the references and follow the (optional) supplemental material.
-The checklist does NOT count towards the page limit.
-Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
-  • You should answer [Yes] , [No] , or [N/A] .
-  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
-  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
-The checklist answers are an integral part of your paper submission.
-They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers.
-You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
-The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation.
-While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used").
-In general, answering "[No] " or "[N/A] " is not grounds for rejection.
-While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate.
-All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix.
-If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
-IMPORTANT, please:
-  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
-  • Keep the checklist subsection headings, questions/answers and guidelines below.
-  • Do not modify the questions and only use the provided macros for your answers.
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4.
-Justification: Yes, the abstract and introduction clearly describe the main contributions, including the design of QGAT and its performance improvements over classical baselines.
-These claims are supported by both the model description and experimental results.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9.
-Justification: The paper includes a discussion of QGAT’s limitations, particularly in terms of computational cost and scalability due to quantum circuit simulation, as well as the restricted hyperparameter space caused by GPU memory constraints.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14.
-Justification: The paper does not present formal theoretical results that require assumptions or proofs.
-The contributions are primarily empirical and architectural, focusing on model design, implementation, and evaluation.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19.
-Justification: The paper provides detailed architectural descriptions, training hyperparameters, and experimental configurations for all tasks.
-This includes dataset-specific training details, model variants, optimizer and scheduler choices, and evaluation metrics.
-These details are sufficient to reproduce the main experimental results, even without code release.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [No]
-  24. Justification: The code is currently not released due to ongoing downstream application development. We plan to make it available after completing further internal experiments.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29.
-Justification: The paper specifies all key training and evaluation settings, including dataset splits, optimizer, learning rate schedules, dropout rates, number of layers and attention heads, and other hyperparameters.
-Detailed training configurations for each benchmark task (transductive, inductive, and link prediction) are provided in the Appendix.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34.
-Justification: All experimental results are reported with standard deviation error bars calculated over five independent runs using different random seeds.
-The error bars in plots and tables reflect variability due to training initialization and stochastic optimization.
-We state explicitly that the bars represent 1-sigma (standard deviation) and reference the corresponding figures (e.g., Figure 2(a), Figure 2(b)) and tables (e.g., Table 1, Table 2) in the main text.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39.
-Justification: All experiments were conducted using a single NVIDIA H100 GPU with 80GB memory.
-Training time per run varied by dataset, ranging from approximately 0.5–3 hours for node classification tasks to 3–6 hours for large-scale link prediction tasks.
-Memory usage was managed using mixed-precision training and sampling strategies (e.g., GraphSAINT).
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44.
-Justification: The research adheres to the NeurIPS Code of Ethics.
-It involves no human subjects, personally identifiable information, sensitive content, or dual-use concerns.
-All datasets used are publicly available and widely adopted in graph learning research.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49.
-Justification: This work explores a quantum-enhanced graph attention architecture (QGAT) for graph learning tasks.
-Potential positive societal impacts include improved performance in scientific domains involving complex graph data, such as drug discovery, protein interaction prediction, and molecular property analysis.
-These applications may accelerate biomedical research and material science.
-    However, as with most machine learning models, the method could potentially be misused in settings like surveillance or social network profiling, where enhanced modeling of graph structures could amplify privacy risks.
-While our work focuses on foundational research and does not target specific applications, we acknowledge this dual-use nature and emphasize that ethical deployment requires domain-specific safeguards and oversight.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54.
-Justification: The paper does not involve the release of high-risk models or scraped datasets.
-QGAT is a foundational architecture evaluated on publicly available benchmark datasets (e.g., OGB, PPI), and does not introduce or train generative models, language models, or datasets with potential safety concerns.
-Therefore, specific safeguards are not applicable in this context.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59.
-Justification: All datasets and software frameworks used in this work are publicly available under appropriate open licenses.
-The datasets (e.g., OGB datasets, PPI) are accessed via the Open Graph Benchmark (OGB) platform, which is released under the MIT license.
-The quantum simulation backend is implemented using Pennylane, which is licensed under Apache 2.0, and the classical components are built with PyTorch, licensed under BSD.
-Citations and references are provided for all external assets used.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., $C$-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [N/A]
-  64.
-Justification: This paper does not introduce any new datasets, models, or code assets.
-All experiments are conducted on publicly available benchmarks and existing libraries, with no new data collection or external asset release.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69.
-Justification: This paper does not involve any experiments with human subjects or the use of crowdsourcing platforms.
-All evaluations are conducted on publicly available benchmark datasets using automated computational methods.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74.
-Justification:This work does not involve any research with human subjects or data collected from human participants.
-All evaluations are performed on publicly available benchmark datasets using computational methods without human intervention.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79.
-Justification: This paper does not incorporate LLMs as part of the core methodology, nor are LLMs used in any original or non-standard way that impacts the scientific contributions of the research.
-All LLM usage, if any, was limited to language editing and does not affect the technical content.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.